# Lightweight and Direct Document Relevance Optimization for Generative Information Retrieval

### v1

@@ 6.2. Comparison with Established Baselines
To address RQ2, Table 3 presents a comprehensive comparison of DDRO with baselines on the MS300K dataset. We can observe the followings: (i) The performance of dense retrieval baselines is generally better than that of sparse retrieval baselines, likely because the former uses dense vectors to capture richer semantic information, which is also consistent with the findings (Lu et al., 2021; Ma et al., 2022). (ii) The best-performing dense retrieval baseline, ANCE, outperforms others such as SEAL, NCI, and DSI-QG. This could be due to the fact that these generative retrieval baselines rely solely on maximum likelihood estimation (MLE) to learn relevance, which may not fully capture the relevance patterns. However, ANCE still lags behind models like Ultron, ROGER, and LTGR, which employ more advanced optimization strategies. This highlights the need for fine-grained relevance modeling to enhance generative retrieval ranking performance. (iii) Our proposed DDRO (TU) outperforms these generative retrieval baselines, achieving 15.63% and 8.03% higher R@1 and MRR@10, respectively, compared to the best-performing baseline, ROGER-Ultron. These results demonstrate the effectiveness of our document-level relevance optimization approach. While ROGER-Ultron achieves the highest R@10, DDRO (TU) delivers comparable performance within a more efficient and lightweight framework. Figure 4 further illustrates the performance of DDRO compared to baselines.
@@
(a) R@1 - MS MARCO
(b) MRR@10 - MS MARCO
Figure 4. Comparison of R@1 and MRR@10 metrics for the MS MARCO dataset across various retrieval models.
[…]
The figure presents two plots comparing R@1 and MRR@10 metrics for various retrieval models on the MS MARCO dataset, along with a shared legend for clarity.
(a) R@1 - Natural Questions
(b) MRR@10 - Natural Questions
Figure 5. Comparison of R@1 and MRR@10 metrics for the Natural Questions dataset across various retrieval models.
[…]
The figure presents two plots comparing R@1 and MRR@10 metrics for various retrieval models on the Natural Questions dataset, along with a shared legend for clarity.
@@ 6.4. Robustness Analysis Across Datasets
Comparison to baseline retrieval models. The performance comparison on the NQ320k across different retrieval baselines is as follow: (i) Term-based baselines, such as BM25 and DocT5Query, show lower early ranking performance, but remain competitive at broader levels. (ii) Dense retrieval baselines, including DPR and ANCE, improve early ranking metrics over term-based methods. (iii) Generative retrieval baselines, such as ROGER-Ultron (TU) and LTRGR, perform well across all metrics. The proposed DDRO achieves the highest overall performance. Specifically, DDRO (PQ) surpasses the best-performing baseline, ROGER-Ultron (TU), by 36.27% in R@1 and 23.58% in MRR@10. Figure 5 further illustrates the performance of DDRO compared to baselines.
@@ 7. Conclusion
We introduced DDRO, a novel approach for enhancing GenIR systems by directly aligning docid generation with document-level relevance estimation. This alignment allows GenIR systems to effectively learn to rank and improve retrieval accuracy. We experimented with two types of docid designs and designed a lightweight direct L2R optimization algorithm on top of SFT training. Unlike existing RL-based methods, DDRO simplifies optimization through pairwise ranking, eliminating the need for auxiliary reward modeling and RL fine-tuning. Experiments conducted on benchmark datasets demonstrate that DDRO offers a lightweight optimization process while achieving high performance and demonstrating robustness. However, DDRO has limitations that offer opportunities for improvement. (i) Our evaluation has primarily focused on a subset of the MS MARCO document dataset, leaving its scalability to larger and more diverse corpora, including domain-specific and multilingual datasets, yet to be explored. Addressing this limitation will be crucial to assess DDRO’s adaptability in broader applications. (ii) Potential improvements could include listwise ranking for improved document ordering and overall ranking effectiveness. Additionally, advanced hard negative mining could enhance relevance discrimination, while multi-objective optimization may help balance relevance with efficiency, fairness, and diversity. (iii) A comprehensive comparison with scalable GenIR baselines such as RIPOR and PAG remains an avenue for future research to assess DDRO’s scalability and integration potential in large-scale retrieval scenarios. Future work will aim to address the mentioned limitations to further evaluate DDRO’s effectiveness and applicability in broader contexts.

### v2

@@ 6.2. Comparison with Established Baselines
To address RQ2, Table 3 presents a comprehensive comparison of DDRO with baselines on the MS300K dataset. We can observe the followings: (i) The performance of dense retrieval baselines is generally better than that of sparse retrieval baselines, likely because the former uses dense vectors to capture richer semantic information, which is consistent with earlier findings (Lu et al., 2021; Ma et al., 2022). (ii) The best-performing dense retrieval baseline, ANCE, outperforms others such as SEAL, NCI, and DSI-QG. This could be due to the fact that these generative retrieval baselines rely solely on maximum likelihood estimation (MLE) to learn relevance, which may not fully capture the relevance patterns. However, ANCE lags behind models like Ultron, ROGER, and LTGR, which employ more advanced optimization strategies. This highlights the need for fine-grained relevance modeling to enhance generative retrieval ranking performance. (iii) Our proposed DDRO (TU) outperforms these generative retrieval baselines, achieving 15.63% and 8.03% higher R@1 and MRR@10, respectively, compared to the best-performing baseline, ROGER-Ultron. These results demonstrate the effectiveness of our document-level relevance optimization approach. While ROGER-Ultron achieves the highest R@10, DDRO (TU) delivers comparable performance within a more efficient and lightweight framework.
@@ 6.3. Ablation Study
Effect of KL Constraint Strength $\beta$.
[…]
A line plot showing MRR@10 across different beta values. Moderate beta (0.4) yields the highest performance on MS MARCO and Natural Questions.
Figure 4. Effect of KL constraint strength ($\beta$) on DDRO performance. A moderate value ($\beta=0.4$) yields the best MRR@10, while under- or over-regularization degrades performance.
@@ 6.4. Robustness Analysis Across Datasets
Comparison to baseline retrieval models. The performance comparison on the NQ320k across different retrieval baselines is as follow: (i) Term-based baselines, such as BM25 and DocT5Query, show lower early ranking performance, but remain competitive at broader levels. (ii) Dense retrieval baselines, including DPR and ANCE, improve early ranking metrics over term-based methods. (iii) Generative retrieval baselines, such as ROGER-Ultron (TU) and LTRGR, perform well across all metrics. The proposed DDRO achieves the highest overall performance. Specifically, DDRO (PQ) surpasses the best-performing baseline, ROGER-Ultron (TU), by 36.27% in R@1 and 23.58% in MRR@10.
@@ 7. Conclusion
We introduced DDRO, a novel approach for enhancing GenIR systems by directly aligning docid generation with document-level relevance estimation. This alignment allows GenIR systems to effectively learn to rank and improve accuracy. We experimented with two types of docid designs and designed a lightweight direct L2R optimization algorithm on top of SFT training. Unlike existing RL-based methods, DDRO simplifies optimization through pairwise ranking, eliminating the need for auxiliary reward modeling and RL fine-tuning. Experiments conducted on benchmark datasets demonstrate that DDRO offers a lightweight optimization process while achieving high performance and demonstrating robustness. However, DDRO has limitations that offer opportunities for further improvement. (i) Our evaluation has primarily focused on a subset of the MS MARCO document dataset, leaving its scalability to larger and more diverse corpora, including domain-specific and multilingual datasets, yet to be explored. Addressing this limitation will be crucial to assess DDRO’s adaptability in broader applications. (ii) Our current pairwise ranking formulation relies on binary relevance judgments (relevant vs. non-relevant), which may limit its expressiveness for queries with nuanced or graded relevance levels. Future work may explore listwise ranking objectives and graded relevance supervision to more accurately capture complex retrieval intents. (iii) Potential improvements could include advanced hard negative mining to enhance relevance discrimination and multi-objective optimization to balance relevance with efficiency, fairness, and diversity. (iv) A comprehensive comparison with scalable GenIR baselines such as RIPOR and PAG remains an avenue for future research to assess DDRO’s scalability and integration potential in large-scale retrieval scenarios.
Acknowledgements.
We thank our colleagues at the IRLab for their support and Hansi Zeng at UMass Amherst for helpful discussions and technical guidance. Experiments for this work were supported by the Dutch Research Council (NWO) under project EINF-9550; computations were performed on the Snellius supercomputer (SURF). This research was (partially) supported by Ahold Delhaize, through AIRLab, by the Dutch Research Council (NWO), under project numbers 024.004.022, NWA.1389.20.183, and KICH3.LTP.20.006, and by the European Union’s Horizon Europe program under grant agreement No 101070212. All content represents the opinion of the authors, which is not necessarily shared or endorsed by their respective employers and/or sponsors.